# Continuum excitations in a spin-supersolid on a triangular lattice

### v2

@@ Abstract
Our first finding is that the magnetic state above 1 T is not a fully polarized one, but an $m=1/3$ magnetization plateau. This is obvious from the magnetization curves measured at $T=1.3$ K on a 5.4 mg single-crystal sample by the coaxial pick-up coil pulsed-field magnetometer at HLD-EMFL [19], and plotted in Fig. 2a. As in all experiments reported below the field is applied along the crystallographic $c$-axis (magnetic easy axis). Note that full saturation is reached only at $\mu_{0}H_{\text{sat}}\approx 22$ T. The difference between the up-sweep and down-sweep curves is instrumental and due to magnetocaloric effect [20]. The same plot shows Faraday balance magnetometry data [21] collected at $190$ mK. The low-temperature value for the start of the plateau phase is $\mu_{0}H_{c}\approx 0.8(1)$ T. Assuming the nearest-neighbor XXZ model, we can link $H_{\text{sat}}$ to the exchange constants for the in- and out-of-plane spin components: $g\mu_{B}\mu_{0}H_{\text{sat}}=3S(J_{xy}+2J_{zz})$, an exact SWT result. From the theoretical ground state phase diagram of the triangular-lattice XXZ model [2, 4], we can thereby estimate $J_{zz}\approx 3.1(3)$ meV and $\alpha=J_{xy}/J_{zz}=0.08(1)$.
A more detailed phase diagram of K2Co(SeO3)2 was established in calorimetric experiments on a Quantum Design PPMS with a dilution refrigerator insert using a 0.05 mg single-crystal sample. The result is presented in a $C/T$ specific heat plot in Fig. 2b [20]. The sharp high-temperature lambda-anomaly reported in [16] is also visible in our data (dotted line). This phase boundary has a well-defined endpoint at ($\mu_{0}H_{0}\approx 1.05$ T, $T_{0}\approx 4.5$ K). At lower field this anomaly is absent. Instead there is a broad crossover line (dashed) that extends down to zero field and corresponds to the bump at $T\sim 1$ K. An additional phase boundary marked by somewhat rounded specific heat maxima is detected at low temperatures. It has the shape of a “dome” of some ordered phase with a zero-field ordering temperature $T_{c}\approx 0.35$ K, and terminates in a quantum critical point $\mu_{0}H_{\text{c}}=0.8$ T, i.e., right at the observed boundary of the magnetization plateau. Overall, the phase diagram is qualitatively similar to that theoretically predicted for Na2BaCo(PO4)2 [5].
@@
The observed magnetic scattering is resolution-limited in the $(h,k)$ plane. Fig. 3b show $l$-scans averaged over two sets of equivalent Bragg reflections [20]. The very broad peaks indicate only short-range correlations along $c$, of approximately AF character. The scans were fit to pairs of Voigt functions centered at $l=\pm 1/2$ (solid lines in Fig. 3b), the Gaussian component fixed at the experimental resolution [20]. From the Lorentzian width we extract the inter-plane correlation length $\lambda\approx 11(1)$, or just 0.6 lattice units. At $\mu_{0}H=1.5$ T the correlation length is only slightly larger [20]. A more detailed analysis is hardly possible given the narrow $l$-coverage.
Figure 3: Elastic scattering measured in K2Co(SeO3)2 in zero field. (a) TOF data at $T=70$ mK in the $(h,k,0)$ plane reveal $(1/3,1/3)$-type two-dimensional order (arrows). Circles are structural Bragg peaks. Solid lines are 2D Brillouin zone boundaries for a single triangular plane. “A” and “B” are two sets of equivalent Bragg points. (b) Background-subtracterd $l$-scans averaged over each set. Solid lines are fits as described in the text. The shaded Gaussian represents experimental resolution as measured on the (-2,1,0) Bragg peak. (c) Temperature dependence of diffraction intensity at $(1/3,1/3,0)$ measured on ZEBRA. The dashed line is the background level measured just off the $(1/3,1/3)$ position.
Only two sets of equivalent Bragg positions are reachable outside of the strong background region at low-$q$. Since the $uud$ and supersolid phases share the same propagation vector, a determination of the 2D magnetic structures does not appear feasible. To get a rough idea of the ordered moment in zero field, we compared the fitted Lorentzian intensities to those determined in $\mu_{0}H=1.5$ T and assumed the latter to be a $uud$ colinear structure. This procedure yields the estimate 2.6(3) $\mu_{\text{B}}$ per Co2+, substantially reduced compared to 3.95 $\mu_{\text{B}}$ at saturation (Fig. 2a).
@@
Figure 4: False-color plots of inelastic neutron intensity measured in K2Co(SeO3)2 at $T=70$ mK in zero field (a and c), and in $\mu_{0}H=1.5$ T (b, note the ×$\times 6$ scale factor). No background is subtracted. The low-energy scattering below 0.3 meV in (b) is spurious, originating from the sample environment. It is also present in (a) but not visible due to a different color range. Panels (d,e) as well as white lines in (a,b) show simulations based on linear SWT. The red line in (d) is the lower bound of the 2-magnon continuum. (f) Equal-time structure factor obtained by integrating the intensity in the range 0.1-0.9 meV.
@@
The contrast is made even more obvious if we compare the observations with linear SWT calculations. We based these on a simple nearest-neighbor XXZ model. All simulations were performed using the SPIN-W software package [27]. First, we used simplistic Gaussian fits to the data to quantify the measured dispersion relation in the $uud$ phase. The result was then fit to linear SWT to obtain the gyromagnetic ratio $g=8.34(6)$ and exchange constants $\tilde{J}_{zz}=0.68(6)$ meV and $\tilde{J}_{xy}=0.25(1)$ meV. The fit is quite unambiguous, since the exchange parameters are well defined by the bandwidths of the upper and lower branch, respectively. The dispersion computed with these values is shown as white lines in Fig. 4a,b. The corresponding intensity simulations are shown in Fig. 4d,e [20], assuming a Gaussian broadening to match experimental energy resolution.
For $\mu_{0}H=1.5$ T SWT reproduces the experimental data up to 1 meV transfer exceptionally well. It also predicts a 3rd mode at a higher energy, one is not seen in the measurement [20]. The obtained effective SWT values $\tilde{J}_{xy}$ and $\tilde{J}_{zz}$, are hugely renormalized compared to the Hamiltonian parameters $J_{xy}$ and $J_{zz}$ deduced from the magnetization curve. This emphasizes that quantum corrections to linear-SWT remain very strong. Similar renormalization has been observed in the plateau phase of triangular-lattice materials such as Ba3CoSb2O9 [6, 7, 8].
@@
In contrast, in zero field the computed low-energy spectra are drastically different from what is actually observed, and not only in what concerns excitation widths. For spinon continua in Heisenberg spin chains [28], the computed shape of the SWT dispersion qualitatively resembles the lower bound of the continuum. Quite to the opposite, in K2Co(SeO3)2 the latter has a distinct dip at the $M$-point, similar to the “roton-like minima” in the easy-plane case [9]. In K2Co(SeO3)2, near the $M$-point, the lower bound of the continuum shows a “relativistic” dispersion with a mass gap $\Delta_{M}=0.12$ meV. This is particulary clear from the constant-energy cut shown Fig. 4c. Weak replicas of the scattering at the $K$ and $M$ points are seen throughout the zone, revealing a hidden $\sqrt{12}\times\sqrt{12}$ periodicity in the dynamic correlations. The large $M$-point intensity contributes to the equal-time structure factor (Fig. 4f) being more or less evenly distributed along the zone boundary, in agreement with numerical simulations for the XXZ model [29].
@@
One possible explanation for the observed behavior is spin wave instability [30], which is allowed in the non-colinear “Y”- structure. This paradigm applies well to systems with rather isotropic interactions, such as Ba3MnSb2O9 [31] or Na2BaCo(PO4)2 [32]. In K2Co(SeO3)2 it is able to account for at least some observed features. In zero field the lower bound of the two-magnon continuum in SWT coincides with the lowest-energy magnon mode (red line in Fig 4d). Above that line all modes other than the lowest energy branch around the $K$ and $\Gamma$ points are prone to decay. Experimentally, this is where the scattering is particularly broad. Even near the $K$ and $\Gamma$ points, the decay of the higher-energy acoustic mode is allowed and accounts for the continuum seen above a sharp lower bound. Unfortunately, this picture provides no explanation for the energy minimum seen at the $M$-point.

### v3

@@ Abstract
Our first finding is that the magnetic state above 1 T is not a fully polarized one, but an $m=1/3$ magnetization plateau. This is obvious from the magnetization curves measured at $T=1.3$ K on a 5.4 mg single-crystal sample by the coaxial pick-up coil pulsed-field magnetometer at HLD-EMFL [19], and plotted in Fig. 2a. As in all experiments reported below the field is applied along the crystallographic $c$-axis (magnetic easy axis). Note that full saturation is reached only at $\mu_{0}H_{\text{sat}}\approx 22$ T. The difference between the up-sweep and down-sweep curves is instrumental and due to magnetocaloric effect [20]. The same plot shows Faraday balance magnetometry data [21] collected at $190$ mK. The low-temperature value for the start of the plateau phase is $\mu_{0}H_{c}\approx 0.8(1)$ T. Assuming the nearest-neighbor XXZ model, we can link $H_{\text{sat}}$ to the exchange constants for the in- and out-of-plane spin components: $g\mu_{B}\mu_{0}H_{\text{sat}}=3S(J_{xy}+2J_{zz})$, an exact SWT result. Previous theoretical work relates the ratio $\alpha\equiv J_{xy}/J_{zz}$ to $H_{\text{sat}}/H_{c}$ [2, 4]. This allows us to estimate $J_{zz}=3.1(3)$ meV and $\alpha=0.08(1)$ from the two measured transition field values.
A more detailed phase diagram of K2Co(SeO3)2 was established in calorimetric experiments on a Quantum Design PPMS with a dilution refrigerator insert using a 0.05 mg single-crystal sample. The result is presented in a $C/T$ specific heat plot in Fig. 2b [20]. The sharp high-temperature lambda-anomaly reported in [16] is also visible in our data (dotted line). This phase boundary has a well-defined endpoint at ($\mu_{0}H_{0}\approx 1.05$ T, $T_{0}\approx 4.5$ K). At lower field this anomaly is absent. Instead there is a broad crossover line (dashed) that extends down to zero field and corresponds to the bump at $T\sim 1$ K. An additional phase boundary marked by somewhat rounded specific heat maxima is detected at low temperatures. It has the shape of a “dome” of some ordered phase with a zero-field ordering temperature $T_{c}\approx 0.35$ K, and terminates in a quantum critical point $\mu_{0}H_{\text{c}}=0.8$ T, i.e., right at the observed boundary of the magnetization plateau. Overall, the phase diagram is qualitatively similar to that theoretically predicted for Na2BaCo(PO4)2, where the sharp lambda-anomaly in field is a Potts transition to the uud phase, replaced by broad features due to BKT transitions at low fields [5].
@@
The observed magnetic scattering is resolution-limited in the $(h,k)$ plane, indicating long-range order in both cases. Fig. 3b show $l$-scans averaged over two sets of equivalent Bragg reflections [20]. The very broad peaks indicate only short-range correlations along $c$, of approximately AF character. The scans were fit to pairs of Voigt functions centered at $l=\pm 1/2$ (solid lines in Fig. 3b), the Gaussian component fixed at the experimental resolution [20]. From the Lorentzian width we extract the inter-plane correlation length $\lambda\approx 11(1)$, or just 0.6 lattice units. At $\mu_{0}H=1.5$ T the correlation length is only slightly larger [20]. A more detailed analysis is hardly possible given the narrow $l$-coverage.
Figure 3: Elastic scattering measured in K2Co(SeO3)2 in zero field. (a) TOF data at $T=70$ mK in the $(h,k,0)$ plane reveal $(1/3,1/3)$-type two-dimensional order (arrows). Circles are structural Bragg peaks. Solid lines are 2D Brillouin zone boundaries for a single triangular plane. “A” and “B” are two sets of equivalent Bragg points that differ only by the absolute value of scattering vector. (b) Background-subtracterd $l$-scans averaged over each set. Solid lines are fits as described in the text. The shaded Gaussian represents experimental resolution as measured on the (-2,1,0) Bragg peak. (c) Temperature dependence of diffraction intensity at $(1/3,1/3,0)$ measured on ZEBRA. The dashed line is the background level measured just off the $(1/3,1/3)$ position.
Only two sets of equivalent Bragg positions are reachable outside of the strong background region at low-$q$. The $uud$ and supersolid phases share the same propagation vector, a determination of the 2D magnetic structures does not appear feasible. To get a rough idea of the ordered moment in zero field, we compared the fitted Lorentzian intensities to those determined in $\mu_{0}H=1.5$ T and assumed the latter to be a $uud$ colinear structure. This procedure yields the estimate 2.6(3) $\mu_{\text{B}}$ per Co2+, substantially reduced compared to 3.95 $\mu_{\text{B}}$ at saturation (Fig. 2a).
@@
Figure 4: False-color plots of inelastic neutron intensity measured in K2Co(SeO3)2 at $T=70$ mK in zero field (a and c), and in $\mu_{0}H=1.5$ T (b, note the ×$\times 6$ scale factor). No background is subtracted. The low-energy scattering below 0.3 meV in (b) is spurious, originating from the sample environment. It is also present in (a) but not visible due to a different color range. The dashed horizontal line in (a) is the position of the constant-energy cut in (c). Panels (d,e) as well as white lines in (a,b) show simulations based on linear SWT. (f) Equal-time structure factor obtained by integrating the intensity in the range 0.1-0.9 meV, chosen to avoid elastic line and any background at higher energies where no obvious signal originating from the lower spectral branches is present.
@@
To emphasize the contrast, we compare the observations with linear SWT calculations. We based these on a simple nearest-neighbor XXZ model with the exchange parameters as quoted above. All simulations were performed using the SPIN-W software package [27]. For the zero-field case, we assumed a tiny $H=30$ G field along the $c$ axis to overcome the accidental classical degeneracy and stabilize the assumed $Y$-structure [20]. For calculations in applied fields we estimated the gyromagnetic ratio $g=7.9$ from the saturation magnetization. The resulting dispersion is shown as white lines in Fig. 4a,b. The corresponding intensity simulations are shown in Fig. 4d,e [20], assuming a Gaussian broadening to match experimental energy resolution.
For $\mu_{0}H=1.5$ T SWT reproduces the data up to 1 meV transfer rather well. A small discrepancy remains in the curvature of the higher-energy branch (see Supplement for a more quantitative discussion). Evidently, quantum corrections to linear-SWT remain significant. Similar effects were observed in the plateau phases of triangular-lattice materials such as Ba3CoSb2O9 [6, 7, 8].
@@
In contrast, in zero field the computed low-energy spectra are drastically different from what is actually observed, and not only in what concerns excitation widths. For spinon continua in Heisenberg spin chains [28], the computed shape of the SWT dispersion qualitatively resembles the lower bound of the continuum. Quite to the opposite, in K2Co(SeO3)2 the latter has a distinct dip at the $M$-point, similar to the “roton-like minima” in the easy-plane case [9]. In K2Co(SeO3)2, near the $M$-point, the lower bound of the continuum shows a “relativistic” dispersion with a mass gap $\Delta_{M}=0.12$ meV. This is particulary clear from the constant-energy cut shown Fig. 4c. Weak replicas of the scattering at the $K$ and $M$ points are seen throughout the zone, revealing a hidden $\sqrt{12}\times\sqrt{12}$ periodicity in the dynamic correlations. The large $M$-point intensity contributes to the equal-time structure factor (Fig. 4f) being more or less evenly distributed along the zone boundary. Overall, our measured low-energy spectra are in excellent agreement with recent numerical simulations for K2Co(SeO3)2 [29, 30].
@@
One possible explanation for the observed behavior is spin wave instability [31], which is allowed in the non-colinear “Y”- structure. This paradigm applies well to systems with rather isotropic interactions, such as Ba3MnSb2O9 [32] or Na2BaCo(PO4)2 [33]. In K2Co(SeO3)2 it is able to account for at least some observed features. In zero field the lower bound of the two-magnon continuum in SWT coincides with the lowest-energy magnon mode. The higher-energy spin wave branch is thus always unstable towards two-particle decay and my give rise to a continuum. Unfortunately, this provides no explanation for the energy minimum seen at the $M$-point. Moreover, the SWT language may be totally inadequate for small $\alpha$. In this limit theory predicts only a tiny transverse order parameter [34]. Acoustic spin waves are the corresponding Goldstone modes, and are nothing other than its precessions. Therefore, they can’t be expected to contribute significantly to the spectrum.